# Human3R: Incorporating Human Priors for Better 3D Dynamic Reconstruction from Monocular Videos

### v1

@@ Abstract
Monocular dynamic video reconstruction faces significant challenges in dynamic human scenes due to geometric inconsistencies and resolution degradation issues. Existing methods lack 3D human structural understanding, producing geometrically inconsistent results with distorted limb proportions and unnatural human-object fusion, while memory-constrained downsampling causes human boundary drift toward background geometry. To address these limitations, we propose to incorporate hybrid geometric priors that combine SMPL human body models with monocular depth estimation. Our approach leverages structured human priors to maintain surface consistency while capturing fine-grained geometric details in human regions. We introduce Human3R, featuring a hierarchical pipeline with refinement components that processes full-resolution images for overall scene geometry, then applies strategic cropping and cross-attention fusion for human-specific detail enhancement. The method integrates SMPL priors through a Feature Fusion Module to ensure geometrically plausible reconstruction while preserving fine-grained human boundaries. Extensive experiments on TUM Dynamics and GTA-IM datasets demonstrate superior performance in dynamic human reconstruction.
Figure 1: Human3R for Dynamic Scene Reconstruction. Our task is to reconstruct accurate 3D point clouds from monocular videos containing humans. By incorporating SMPL human priors, Human3R achieves superior reconstruction quality for both dynamic human motions and static scene elements compared to existing methods. The highlighted regions demonstrate our method’s ability to preserve anatomically consistent human geometry and fine-grained details while maintaining overall scene coherence.
@@ 1 Introduction
To address the aforementioned challenges, we propose a novel method, Human3R. To mitigate the Anatomical Artifacts issue, we introduce a hybrid geometric prior that combines the Skinned Multi-Person Linear (SMPL) model with monocular depth estimation. Specifically, the SMPL model provides a consistent human surface prior, while monocular depth estimation contributes fine-grained geometric details. By aligning the SMPL geometry with the monocular depth map, we obtain detailed depth predictions that maintain strong surface consistency. Furthermore, to tackle the Resolution and Boundary Degradation issue, we design a pipeline with a refinement component. The network initially processes the full-resolution image to capture the overall scene geometry. When humans appear small in the image, our refinement module applies strategic image cropping to effectively super-resolve human regions. A dedicated branch then employs cross-attention between the enlarged human region and the original image context, enabling precise reconstruction of human-like forms and ultimately producing more detailed and accurate human reconstruction results.
Figure 2: The overview of our framework. The left shows the base Human3R pipeline that processes image pairs through DUSt3R encoders, followed by a Feature Fusion Module that integrates image features, monocular depth features, and SMPL depth features via cross-attention mechanism and gated operations. The fused features are then processed through transformer decoders to produce initial point map reconstruction. The right illustrates Human3R with Refinement, which is activated when humans occupy small pixel regions in the input images. The refinement stage crops human regions from the input images, upsamples the point clouds, and processes them through refinement decoders that incorporate cross-module attention between local crop features and preserved scene context. This hierarchical design enables high-fidelity human reconstruction while maintaining geometric consistency between human regions and the surrounding scene.
@@
  • We propose a novel model, Human3R, capable of accurate dynamic human reconstruction.
@@ 3.1 Overview
Given a video consisting of $N$ frames $\{\mathbf{I}_{k}\in\mathbb{R}^{H\times W\times 3}\}_{k=1}^{N}$, our target is to estimate the corresponding depth maps $\{\mathbf{D}_{k}\in\mathbb{R}^{H\times W}\}_{k=1}^{N}$ and camera poses $\{\pi_{k}\in\mathbb{SE}(3)\}_{k=1}^{N}$. Human3R achieves this through a hierarchical pipeline with a refinement:
@@ 4.1 Experimential Setting
Our Human3R is initialized with the weight of Align3R [lu2024align3ralignedmonoculardepth] and trained using AdamW [Loshchilov2017DecoupledWD] with a learning rate of $5\times 10^{-5}$ and a batch size of 5 on 7 NVIDIA H800 GPUs. Training our model on $5\times 10^{4}$ pairs for 50 epochs in each of main pipeline and refinement separately takes approximately one week in total.
@@ 4.2 Video Depth Estimation
We compare our Human3R with 1 diffusion-based 4D feed forward reconstruction method: Geo4D [jiang2025geo4dleveragingvideogenerators], and three non-diffusion-based methods: MonST3R [zhang2025monst3rsimpleapproachestimating], Align3R [lu2024align3ralignedmonoculardepth], and VGGT [wang2025vggtvisualgeometrygrounded].
@@ Results.
Table 1 presents the quantitative results for video depth estimation. Our Human3R demonstrates superior performance on real-world datasets, achieving the best results on BEHAVE (Abs Rel: 0.033, $\delta<1.25$: 0.992), Bonn (Abs Rel: 0.077), and TUM Dynamics (Abs Rel: 0.102, $\delta<1.25$: 0.907). On synthetic datasets, our method shows competitive performance on GTA-IM (second-best Abs Rel: 0.112, best $\delta<1.25$: 0.869), while Geo4D performs better on Bedlam due to its diffusion-based nature. Overall, our approach effectively handles human-centric scenes with complex interactions.
@@
Table 2 shows the camera pose estimation results. Our Human3R achieves the best performance on synthetic GTA-IM dataset across all metrics (ATE: 0.015, RTE: 0.010, RRE: 0.227), demonstrating significant improvements in temporal consistency and geometric coherence for human-centric scenes. On real-world datasets, our method shows competitive performance, achieving second-best ATE on TUM Dynamics (0.335) and second-best RTE on Bonn (0.285), while obtaining the best RRE on TUM Dynamics (0.430).
@@ 4.4.1 Prior Incorporation
Figure 4: Qualitative comparison of human reconstruction on Human3R with and without refinement. Our refinement module produces more accurate human geometry and maintains better temporal consistency across frames in the cropped human regions.
@@ 5 Conclusion
We proposed Human3R, a novel approach for monocular dynamic video reconstruction in human-centric scenes that addresses anatomical artifacts and resolution degradation issues. Our method incorporates hybrid geometric priors combining SMPL human body models with monocular depth estimation for anatomically plausible reconstruction.
@@ 8.1 Camera pose comparison
Figure 7 demonstrates the camera pose estimation performance of our Human3R method compared to existing approaches on the GTA-IM dataset. Our method achieves superior trajectory accuracy and temporal consistency, particularly in human-centric scenes.

### v2

# HuPrior3R: Incorporating Human Priors for Better 3D Dynamic Reconstruction from Monocular Videos

@@ Abstract
Monocular dynamic video reconstruction faces significant challenges in dynamic human scenes due to geometric inconsistencies and resolution degradation issues. Existing methods lack 3D human structural understanding, producing geometrically inconsistent results with distorted limb proportions and unnatural human-object fusion, while memory-constrained downsampling causes human boundary drift toward background geometry. To address these limitations, we propose to incorporate hybrid geometric priors that combine SMPL human body models with monocular depth estimation. Our approach leverages structured human priors to maintain surface consistency while capturing fine-grained geometric details in human regions. We introduce HuPrior3R, featuring a hierarchical pipeline with refinement components that processes full-resolution images for overall scene geometry, then applies strategic cropping and cross-attention fusion for human-specific detail enhancement. The method integrates SMPL priors through a Feature Fusion Module to ensure geometrically plausible reconstruction while preserving fine-grained human boundaries. Extensive experiments on TUM Dynamics and GTA-IM datasets demonstrate superior performance in dynamic human reconstruction.
Figure 1: HuPrior3R for Dynamic Scene Reconstruction. Our task is to reconstruct accurate 3D point clouds from monocular videos containing humans. By incorporating SMPL human priors, HuPrior3R achieves superior reconstruction quality for both dynamic human motions and static scene elements compared to existing methods. The highlighted regions demonstrate our method’s ability to preserve anatomically consistent human geometry and fine-grained details while maintaining overall scene coherence.
@@ 1 Introduction
To address the aforementioned challenges, we propose a novel method, HuPrior3R. To mitigate the Anatomical Artifacts issue, we introduce a hybrid geometric prior that combines the Skinned Multi-Person Linear (SMPL) model with monocular depth estimation. Specifically, the SMPL model provides a consistent human surface prior, while monocular depth estimation contributes fine-grained geometric details. By aligning the SMPL geometry with the monocular depth map, we obtain detailed depth predictions that maintain strong surface consistency. Furthermore, to tackle the Resolution and Boundary Degradation issue, we design a pipeline with a refinement component. The network initially processes the full-resolution image to capture the overall scene geometry. When humans appear small in the image, our refinement module applies strategic image cropping to effectively super-resolve human regions. A dedicated branch then employs cross-attention between the enlarged human region and the original image context, enabling precise reconstruction of human-like forms and ultimately producing more detailed and accurate human reconstruction results.
Figure 2: The overview of our framework. The left shows the base HuPrior3R pipeline that processes image pairs through DUSt3R encoders, followed by a Feature Fusion Module that integrates image features, monocular depth features, and SMPL depth features via cross-attention mechanism and gated operations. The fused features are then processed through transformer decoders to produce initial point map reconstruction. The right illustrates HuPrior3R with Refinement, which is activated when humans occupy small pixel regions in the input images. The refinement stage crops human regions from the input images, upsamples the point clouds, and processes them through refinement decoders that incorporate cross-module attention between local crop features and preserved scene context. This hierarchical design enables high-fidelity human reconstruction while maintaining geometric consistency between human regions and the surrounding scene.
@@
  • We propose a novel model, HuPrior3R, capable of accurate dynamic human reconstruction.
@@ 3.1 Overview
Given a video consisting of $N$ frames $\{\mathbf{I}_{k}\in\mathbb{R}^{H\times W\times 3}\}_{k=1}^{N}$, our target is to estimate the corresponding depth maps $\{\mathbf{D}_{k}\in\mathbb{R}^{H\times W}\}_{k=1}^{N}$ and camera poses $\{\pi_{k}\in\mathbb{SE}(3)\}_{k=1}^{N}$. HuPrior3R achieves this through a hierarchical pipeline with a refinement:
@@ 4.1 Experimential Setting
Our HuPrior3R is initialized with the weight of Align3R [lu2024align3ralignedmonoculardepth] and trained using AdamW [Loshchilov2017DecoupledWD] with a learning rate of $5\times 10^{-5}$ and a batch size of 5 on 7 NVIDIA H800 GPUs. Training our model on $5\times 10^{4}$ pairs for 50 epochs in each of main pipeline and refinement separately takes approximately one week in total.
@@ 4.2 Video Depth Estimation
We compare our HuPrior3R with 1 diffusion-based 4D feed forward reconstruction method: Geo4D [jiang2025geo4dleveragingvideogenerators], and three non-diffusion-based methods: MonST3R [zhang2025monst3rsimpleapproachestimating], Align3R [lu2024align3ralignedmonoculardepth], and VGGT [wang2025vggtvisualgeometrygrounded].
@@ Results.
Table 1 presents the quantitative results for video depth estimation. Our HuPrior3R demonstrates superior performance on real-world datasets, achieving the best results on BEHAVE (Abs Rel: 0.033, $\delta<1.25$: 0.992), Bonn (Abs Rel: 0.077), and TUM Dynamics (Abs Rel: 0.102, $\delta<1.25$: 0.907). On synthetic datasets, our method shows competitive performance on GTA-IM (second-best Abs Rel: 0.112, best $\delta<1.25$: 0.869), while Geo4D performs better on Bedlam due to its diffusion-based nature. Overall, our approach effectively handles human-centric scenes with complex interactions.
@@
Table 2 shows the camera pose estimation results. Our HuPrior3R achieves the best performance on synthetic GTA-IM dataset across all metrics (ATE: 0.015, RTE: 0.010, RRE: 0.227), demonstrating significant improvements in temporal consistency and geometric coherence for human-centric scenes. On real-world datasets, our method shows competitive performance, achieving second-best ATE on TUM Dynamics (0.335) and second-best RTE on Bonn (0.285), while obtaining the best RRE on TUM Dynamics (0.430).
@@ 4.4.1 Prior Incorporation
Figure 4: Qualitative comparison of human reconstruction on HuPrior3R with and without refinement. Our refinement module produces more accurate human geometry and maintains better temporal consistency across frames in the cropped human regions.
@@ 5 Conclusion
We proposed HuPrior3R, a novel approach for monocular dynamic video reconstruction in human-centric scenes that addresses anatomical artifacts and resolution degradation issues. Our method incorporates hybrid geometric priors combining SMPL human body models with monocular depth estimation for anatomically plausible reconstruction.
@@ 8.1 Camera pose comparison
Figure 7 demonstrates the camera pose estimation performance of our HuPrior3R method compared to existing approaches on the GTA-IM dataset. Our method achieves superior trajectory accuracy and temporal consistency, particularly in human-centric scenes.